# DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models

### v2

@@ 2.3 Data Heterogeneity
The second challenge is the computational imbalance stemming from data heterogeneity. Each input sequence (i.e., training sample) for multimodal LLM training consists of interleaved modality subsequences that exhibit highly skewed distributions. Focusing on images and texts, we perform data characterization on the LAION-400M dataset [30], an open-source collection of images paired with text captions. Each image (i.e., one image subsequence) is segmented into 1616 patches, and each patch is tokenized into one image token. The texts are tokenized through Llama tokenizer. The image tokens are interleaved with text tokens to create an 8K-token input sequence for training. As shown in Figure 5(a) and Figure 5(b), the sizes of text and image subsequences display distinctly skewed distributions. We further analyze the count of modality subsequence per training sample using image as an example. The count of image subsequences per training sample, shown in Figure 5(c), also demonstrates a skewed distribution. Different sample size (i.e., modality tokens per sample) leads to varying computation time in the modality encoder and generator stages.
@@
Such data heterogeneity results in both intra-microbatch and inter-microbatch stragglers within the pipeline parallelism (PP) stages of the modality encoder and generator. These stragglers exacerbate the computational imbalances and further reduce the GPU utilization. It is noted that all microbatches within LLM backbone have the same computation time since the sequence length is fixed. We do not consider data heterogeneity between global batches, as each global batch contains numerous randomly shuffled training samples (e.g., thousands with a large DP size), which effectively smooths out the data heterogeneity.
@@
Intra-microbatch straggler. Intra-microbatch straggler occurs when particularly large training samples decelerate training, as DP groups handle variably-sized training samples. Illustrated in Figure 6, the first DP group () processes two large training samples within two microbatches. In contrast, the second DP group () processes two smaller samples in the same microbatches, completing them more swiftly. Consequently, lags behind and becomes the straggler, which delays the overall training process.
@@
Inter-microbatch straggler. Inter-microbatch straggler emerges from pipeline imbalances between microbatches. As depicted in Figure 7, the first pipeline stage is the modality encoder followed by one LLM backbone stage. Figure 7(a) illustrates the pipeline without data heterogeneity, where the modality encoder processes each microbatch with consistent time. In contrast, Figure 7(b) depicts the pipeline with data heterogeneity, where the forward time of the modality encoder varies markedly across microbatches. The straggler (i.e., the microbatch ) significantly delays the training process of the subsequent PP stages, leading to a large pipeline bubble.
3 DistTrain Overview
We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity. DistTrain proposes disaggregated training for multimodal LLM to achieve high efficiency and scalability. DistTrain eliminates the model heterogeneity by disaggregated model orchestration (§5) and harnesses the data heterogeneity by disaggregated data reordering (§6). In addition, DistTrain adopts some system optimizations tailored for multimodal LLM training (§7). Here we provide a brief overview of DistTrain as Figure 8 shows.
@@ 3 DistTrain Overview
DistTrain manager. Before training, DistTrain employs a training manager to determine the resource allocation and parallelism strategy for each module in multimodal LLMs. This scheduler first gathers the model architecture and training configuration (e.g., global batch size) from the user and randomly samples a subset of training data to analyze the data distribution. Utilizing this information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time. Based on the profiling results, the training manager decides the optimal resource allocation and parallelism strategy with disaggregated model orchestration for one specific training task, as detailed in §5.
@@
DistTrain initializer. DistTrain then initializes the GPU training disaggregation for modality encoder, LLM backbone, and modality generator, respectively. DistTrain allocates different numbers of GPUs to each parallelism unit. Each unit then establishes the specific communication group. The unit loads the model checkpoint from distributed file system and shards the model parameters and optimization states. Finally, DistTrain conducts several communication trials to warm up the system and test connectivity.
@@
DistTrain runtime. At runtime, the dedicated CPU nodes (i.e., CPU preprocessing disaggregation) retrieve training samples from the distributed file system for preprocessing. It performs disaggregated data reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31]. This reordering effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §6. In each iteration, the main training process receives the preprocessed data asynchronously from the CPU nodes. The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline. Finally, it synchronizes the gradients and model parameters through the all-gather operation, employing the ZERO-1 optimization [32] and mixed precision training [33]. Additionally, DistTrain adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
4 Disaggregated Training
To address the model and data heterogeneity in multimodal LLM training, we first introduce DistTrain’s core principle: disaggregated training. It includes GPU training disaggregation and CPU preprocessing disaggregation. GPU training disaggregation provides opportunities for adaptive model orchestration across the three modules to address the model heterogeneity. CPU preprocessing disaggregation facilitates data preprocessing to address data heterogeneity with negligible runtime overhead.
4.1 GPU Training Disaggregation
Figure 9 demonstrates the training topology of disaggregated GPU training in DistTrain. Different from the rigid model orchestration in Megatron-LM (i.e., Figure 2), DistTrain is able to adaptively adjust the resource allocation and parallelism strategy. For instance, DistTrain allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator. Additionally, the projector layers are co-located with either the modality encoder or generator, with their number of replicas adapting as needed. We implement GPU training disaggregation through a dedicated module, i.e., parallelism unit.
Parallelism unit. At training initialization, we need to establish the communication group according to the resource allocation and parallelism strategy. DistTrain introduces a module, parallelism unit, composed of one or more PP stages. Each unit can adopt its own DP and TP strategies and form a specific communication group. Inter-unit connections are facilitated by a communication broker, which bridges PP communication across parallelism units. Users are only required to specify the DP and TP configurations for each parallelism unit, and DistTrain automatically sets up the communication group and communication broker. DistTrain treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units. The detailed implementation of parallelism unit is discussed in §7.
4.2 CPU Preprocessing Disaggregation
When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data. The latter significantly increases data preprocessing time. For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images. The text is just kilobytes, whereas the images are total of 120 megabytes. Preprocessing (e.g., decompression, resizing, and reordering) such samples can take several seconds and interfere with the co-located training process. DistTrain disaggregates the CPU data preprocessing from the GPU training process with a producer-consumer model. The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and preprocesses training data asynchronously with the GPU training process. The consumer, i.e., the main training process, receives this preprocessed data for training. The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available. This disaggregation guarantees that the CPU preprocessing does not interfere with the GPU training process and achieves negligible data preprocessing overhead.
5 Addressing Model Heterogeneity
@@
Disaggregated training enables disaggregated model orchestration among the different modules based on the training workload. We first formulate the problem of disaggregated model orchestration to minimize the training time per iteration. Then, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
5.1 Problem Formulation
@@
With disaggregated training, we are able to adaptively orchestrate the three modules. The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration. Exhaust search is infeasible due to the large search space, particularly in large cluster. One strawman solution is to allocate the resources proportional to the model flops of each module. However, this method falls short as it overlooks complex patterns in parallelism training. Before diving into disaggregated model orchestration, we first formulate the optimization problem.
@@
LLM backbone. We begin by formulating the LLM backbone, focusing on the forward pass as the backward pass time mirrors this. In LLM training, microbatch size is set to one to prevent GPU memory overflow. Assume the global batch size for one iteration as and the TP size of the LLM backbone as . Let the PP and DP size of LLM backbone be and . The number of GPUs allocated to LLM backbone is . Let the forward time (including communication time) of the entire LLM be , where represents forward time function. Therefore, the forward time of one PP stage for one microbatch is . Besides, the number of microbatches per iteration is .
@@
Modality encoder and generator. In DistTrain, the modality encoder is regarded as a parallelism unit with PP size . Let the TP size be and the DP size be . The number of GPUs allocated to modality encoder is . The microbatch size is which is determined by the LLM backbone. Let the forward time (including communication time) of the entire modality encoder be . The forward time of one PP stage for one microbatch in the modality encoder is . Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality generator.
@@
Objective function. Based on the preceding analysis, we next define the objective function for the optimization problem, i.e., the training time of one iteration. As shown in Figure 10, we demonstrate the pipeline of forward pass in multimodal LLM training. The LLM backbone comprises two PP stages, whereas the modality encoder and generator each consist of one PP stage. The training process is categorized into two phases: warm-up phase and steady phase. The warm-up phase spans from the initiation to the completion of the first microbatch to populate the pipeline. This phase’s duration is calculated as , which is formulated as follows.
@@
The steady phase’s duration is dominated by the maximal computation time among PP stages, which is calculated as , where is the number of microbatches per iteration. It is formulated as:
@@
Therefore, the objective function is to minimize . For the backward pass, the objective function remains analogous to that of the forward pass. Adjustments are made by changing , and from forward time functions to the sum functions of forward and backward time. This formulation holds for GPipe and 1F1B. We will retrofit the formulation to adapt to VPP later. TP communication is incorporated into the functions , and , which are calibrated through interpolation from actual trials. The communication time of DP and PP is modeled as the communication volume divided by the bandwidth. If the DP sizes are fixed, the DP communication time remains constant. For PP, the communication time equals the PP size multiplied by the communication time of a single layer’s output tensor.
@@
Constraints. Besides the objective function, we must consider constraints to ensure training feasibility. The first constraint is the resource constraint. The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster. The second constraint involves memory. We consider LLM backbone. Memory allocation involves four parts: model parameters, gradients, optimizer states, and activation states. The memory of model parameters and gradients on one GPU is calculated as: , where denotes total memory for the LLM parameters and gradients. The memory for optimizer states on one GPU (with ZeRO-1 optimization) is: , where denotes the total memory for the optimizer states. ZeRO-1 partitions the optimizer states across DP groups. The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch of activation states across the entire LLM. In 1F1B, the first PP stage requires storage for microbatches of activation states. We eschew using GPipe in DistTrain since GPipe consumes more memory. The memory constraint ensures the sum of the four memory parts on one GPU doesn’t exceed GPU capacity. As for modality encoder and generator, the formulation is similar.
5.2 Disaggregated Model Orchestration
@@
The optimization problem is non-convex, with , and the DP, TP sizes as positive variables. is predefined by the user. Solving this with an exhaust search algorithm to explore all possible variable values is impractical due to the extensive search space, particularly in large clusters. Designing an efficient algorithm that quickly identifies the optimal resource allocation and parallelism strategy is a significant challenge.
@@
Convex optimization. Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., resource allocation). We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups. The PP size of the LLM backbone is calculated as . The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of TP and DP size. This allows us to enumerate all feasible TP and DP sizes and transform the original optimization problem into a set of simplified problems. In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive. Therefore, the objective function is convex. Similarly, the constraint functions are also convex. As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by existing solvers [34, 35]. The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy.
@@
Virtual pipeline parallelism (i.e., interleaved 1F1B) reduces the warm-up time by dividing model into finer-grained virtual PP (VPP) stages. Each PP stage contains VPP-size VPP stages. In the warm-up phase, each PP stage launches the computation of one VPP stage, and the warm-up time is divided by VPP-size. To align our formulation with VPP, we proportionally reduce the warm-up time based on VPP-size.
6 Addressing Data Heterogeneity

### v3

@@ 2.3. Data Heterogeneity
Characterization. The training data samples of Multimodal LLMs often combine lightweight text with heavyweight multimodal data. The latter significantly increases data preprocessing time. For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images. The text is just kilobytes, whereas the images are total of 120 megabytes. Preprocessing (e.g., decompression, resizing, and reordering) such samples can take several seconds, and prolong the co-located training process. Additionally, each input sequence consists of interleaved modality subsequences that exhibit highly skewed distributions. Focusing on images and texts, we perform data characterization on the LAION-400M dataset (Schuhmann et al., 2021). Each image (i.e., one image subsequence) is segmented into 1616 patches, and each patch is converted into one image token. The texts are tokenized through Llama tokenizer. The image tokens are interleaved with text tokens to create an 8K-token training sequence. As shown in Figure 5(a) and Figure 5(b), the sizes of text and image subsequences display highly skewed distributions. We further analyze the count of modality subsequence per training sample using image as an example. The count of image subsequences per training sample, shown in Figure 5(c), also shows a skewed distribution. Different sample sizes (i.e., modality tokens per sample) lead to varying computation time in the modality encoder and generator.
@@
Stragglers. Data heterogeneity results in intra-microbatch and inter-microbatch stragglers within the PP stages of the modality encoder and generator. These stragglers exacerbate the computation imbalances and further reduce GPU utilization. It is noted that all microbatches within the LLM have the same computation time since the sequence length is fixed. We do not consider data heterogeneity between global batches, as each global batch contains numerous randomly shuffled training samples (e.g., thousands with a large DP size), which effectively smooths out data heterogeneity.
@@
Intra-microbatch straggler arises since different DP groups handle variably-sized training samples. Illustrated in Figure 6, the first DP group () processes two large training samples within two microbatches. In contrast, the second DP group () processes two smaller samples in the same microbatches, completing them more quickly. Consequently, lags behind and becomes the straggler, which delays the overall training process.
@@
Inter-microbatch straggler emerges from pipeline imbalances between microbatches. As depicted in Figure 7, the first pipeline stage is the modality encoder followed by one LLM backbone stage. Figure 7(a) illustrates the pipeline without data heterogeneity, where the modality encoder processes each microbatch with the same amount of time. In contrast, Figure 7(b) depicts the pipeline with data heterogeneity, where the forward time of the modality encoder varies largely across microbatches. The straggler (i.e., the microbatch a) significantly delays the training process of the subsequent PP stages, leading to a large pipeline bubble.
3. DistTrain Overview
We propose DistTrain, a disaggregated training system for multimodal LLMs. DistTrain addresses model heterogeneity by disaggregated model orchestration (§4) and handles data heterogeneity by disaggregated data preprocessing (§5). Figure 8 shows an overview of DistTrain.
@@ 3. DistTrain Overview
DistTrain manager. Before training, DistTrain employs a training manager to determine the resource allocation and parallelism strategy for each module in multimodal LLMs. The training manager first gathers the model architecture and training configuration (e.g., global batch size) from the user and samples a subset of training data to analyze the data distribution. Utilizing the information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time. Based on the profiling results, the training manager decides the optimal resource allocation and parallelism strategy with disaggregated model orchestration for one specific training task, as detailed in §4.
@@
DistTrain initializer. DistTrain then initializes the disaggregated model orchestration for modality encoder, LLM backbone, and modality generator, respectively. DistTrain allocates different numbers of GPUs and determines the parallelism strategy for each module. Each module then establishes the specific communication group. It then loads the model checkpoint from a distributed file system and shards the model parameters. Finally, DistTrain conducts several communication trials to warm up the system and test connectivity.
@@
DistTrain runtime. At runtime, the dedicated CPU nodes (i.e., disaggregated data preprocessing nodes) retrieve training data samples from the distributed file system for preprocessing. It performs data reordering to reorder the training samples within one global batch while preserving the synchronous training semantics (alp, 2022). The reordering mitigates both intra-microbatch and inter-microbatch stragglers caused by data heterogeneity, as detailed in §5. In each iteration, the GPU training nodes receive the preprocessed data asynchronously from the CPU nodes. The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline. Finally, the GPU training nodes synchronize the gradients and model parameters through collective communication. DistTrain employs ZERO-1 optimization (Rajbhandari et al., 2020) and mixed precision training (Micikevicius et al., 2017) to reduce the memory footprint and accelerate the training of LLM backbone. Additionally, DistTrain adopts a dedicated process to periodically and asynchronously save model checkpoints to the distributed file system for fault tolerance.
4. Addressing Model Heterogeneity
@@
To address model heterogeneity, we first introduce disaggregated model orchestration to adaptively orchestrate the resource and parallelism configurations of the three modules in multimodal LLM training. We then formulate the problem to minimize the training time per iteration. Last, we present the adaptive model orchestration algorithm to find optimal resource and parallelism configurations.
4.1. Disaggregated Model Orchestration
Figure 9 illustrates the disaggregated model orchestration. Different from the monolithic model orchestration in Megatron-LM (i.e., Figure 2), DistTrain is able to adaptively adjust the resource allocation and parallelism strategy. For instance, DistTrain can allocate 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator. Additionally, the projector layers are co-located with either the modality encoder or generator, with their number of replicas adapting as needed. We implement disaggregated model orchestration through a dedicated module, i.e., parallelism unit.
Parallelism unit. At training initialization, we need to establish the communication group according to the resource allocation and parallelism strategy. DistTrain introduces a module, parallelism unit, composed of one or more PP stages. Each unit can adopt its own DP and TP strategies and form a specific communication group. Inter-unit connections are facilitated by a communication broker, which bridges PP communication across parallelism units. Users are only required to specify the DP and TP configurations for each parallelism unit, and DistTrain automatically sets up the communication group and communication broker. DistTrain treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units. For advanced workloads, DistTrain extends the capabilities of these units. To handle long sequences, it integrates sequence parallelism (SP) within the LLM backbone unit and automatically splits sequences within its SP group. Furthermore, DistTrain supports expert parallelism (EP) for the LLM backbone. Since EP and TP both perform parallel computation and communication within one layer, our subsequent formulation involving TP remains valid when TP is replaced with EP. The detailed implementation of parallelism unit is described in §6.
4.2. Problem Formulation
@@
With disaggregated model orchestration, we are able to adaptively orchestrate the three modules. The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration. Exhaust search is infeasible due to the combinatorial search space, particularly for large clusters. One strawman solution is to allocate the resources proportional to the model flops of each module. However, this method falls short as it overlooks complex patterns in parallelism training. Different DP and TP sizes in modality encoder could lead to different computation times while the model flops are the same. Before diving into our solution, we first formulate the optimization problem.
@@
LLM backbone. We begin by formulating the LLM backbone, focusing on the forward pass as the backward pass mirrors it. In production LLM training, the microbatch size is set to a small number to prevent GPU memory overflow. DistTrain sets the microbatch size to a predefined constant, . Assume the global batch size for one iteration as and the TP size of the LLM backbone as . Let the PP and DP size of the LLM backbone be and . The LLM is divided into equal PP stages, where each PP stage consists of some homogeneous transformer layers. The number of GPUs allocated to the LLM backbone is . Let the forward time (including communication time) of the entire LLM for one sample be , where represents the forward time function. Therefore, the forward time of one PP stage for one microbatch with samples is since PP partitions LLM backbone’s homogeneous layers sequentially. Besides, the number of microbatches per iteration is .
@@
Modality encoder and generator. In DistTrain, the modality encoder is regarded as a parallelism unit with PP size . Let the TP size be and the DP size be . The number of GPUs allocated to the modality encoder is . The microbatch size is which is determined by the LLM backbone. Let the forward time of the entire modality encoder for one training sample be . The forward time of one PP stage for one microbatch in the modality encoder is . Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality generator.
@@
Objective function. Based on the preceding analysis, we next define the objective function for the optimization problem, i.e., the training time of one iteration. Figure 10 shows the pipeline of forward pass in multimodal LLM training. The LLM backbone comprises two PP stages, whereas the modality encoder and generator each consist of one PP stage. The forward pass is categorized into two phases: warm-up phase and steady phase. The warm-up phase spans from the initiation to the completion of the first microbatch to populate the pipeline. The duration of this phase is calculated as , which is formulated as follows.
@@
The duration of steady phase is dominated by the maximal computation time among PP stages, which is calculated as , where is the number of microbatches per iteration. It is formulated as:
@@
Therefore, the objective function is to minimize . Considering backward pass with the similar computation pattern, the objective function is similar to that of the forward pass. Adjustments are made by changing , and from forward time functions to the sum functions of forward and backward time. This formulation holds for GPipe (Huang et al., 2019) and 1F1B (Fan et al., 2021). We will retrofit the formulation to adapt to VPP (Narayanan et al., 2021) later. TP communication is incorporated into the functions , and , which are calibrated through interpolation from actual trials. The communication time of DP and PP is modeled as the communication volume divided by the bandwidth.
@@
Constraints. Besides the objective function, we must consider constraints to ensure training feasibility. The first constraint is the resource constraint. The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster and are the number of GPUs allocated to each module in multimodal LLM. The second constraint involves GPU memory. We first consider LLM backbone. Memory allocation involves four parts: model parameters, gradients, optimizer states, and activation states. The memory of model parameters and gradients on one GPU is calculated as: , where denotes the total memory for the LLM parameters and gradients. The memory for optimizer states on one GPU (with ZeRO-1 optimization (Rajbhandari et al., 2020)) is: , where denotes the total memory for the optimizer states. ZeRO-1 partitions the optimizer states across DP groups. The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch (with microbatch size ) of activation states across the entire LLM. In 1F1B, the first PP stage requires storage for microbatches of activation states. We do not use GPipe in DistTrain since it consumes more memory without offering better training efficiency compared to 1F1B. The memory constraint ensures the sum of the four memory parts on one GPU does not exceed GPU capacity. As for the modality encoder and generator, the formulation is similar.
4.3. Adaptive Model Orchestration
@@
The problem is non-convex, with , and the DP, TP sizes as positive variables. Solving this with an exhaust search is impractical due to the large search space, particularly in large clusters. Designing an efficient algorithm that adaptively identifies the optimal resource allocation and parallelism strategy based on the training task poses a hard problem.
@@
Convex optimization. Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., resource allocation). We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups. If expert parallelism (EP) is used instead of TP, we confine EP size to factors of cluster size. The PP size of the LLM backbone is calculated as . The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of TP and DP size. This allows us to enumerate all feasible TP and DP sizes, and transform the original optimization problem into a set of simplified problems. In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive variables. Therefore, the objective function is convex. Similarly, the constraint functions are also convex. As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by off-the-shelf solvers (Grant and Boyd, 2014; Virtanen et al., 2020). The algorithm adaptively determines the optimal resource allocation and parallelism strategy.
@@
Virtual pipeline parallelism (Narayanan et al., 2021) (i.e., interleaved 1F1B) reduces the warm-up time by dividing model into finer-grained virtual PP (VPP) stages. Each PP stage contains VPP-size VPP stages. In the warm-up phase, each PP stage launches the computation of one VPP stage, and the warm-up time is divided by VPP-size. To align our formulation with VPP, we proportionally reduce the warm-up time based on VPP-size.
5. Addressing Data Heterogeneity